Required information

[The following information applies to the questions displayed below.]

Warnerwoods Company uses a perpetual inventory system. It entered into the following purchases and sales transactions for March.

Date Activities Units Acquired at Cost Units Sold at Retail
Mar. 1 Beginning inventory 230 units @ $53.60 per unit
Mar. 5 Purchase 290 units @ $58.60 per unit
Mar. 9 Sales 390 units @ $88.60 per unit
Mar. 18 Purchase 150 units @ $63.60 per unit
Mar. 25 Purchase 280 units @ $65.60 per unit
Mar. 29 Sales 260 units @ $98.60 per unit
Totals 950 units 650 units
Perpetual LIFO:
Goods Purchased Cost of Goods Sold Inventory Balance
Date # of units Cost per unit # of units sold Cost per unit Cost of Goods Sold # of units Cost per unit Inventory Balance
March 1 230 @ $53.60 = $12,328.00
March 5
March 9
March 18
March 25
March 29
Totals $0.00
Answer #1

Under LIFO ,units acquired last are sold first .

Perpetual LIFO
DATE GOODS PURCHASED COST OF GOODS SOLD INVENTORY BALANCE
Units unit cost Total Units unit cost Total Units unit cost Total
Mar 1 230 53.6 12328
Mar 5 290 58.6 16994 230 53.6 12328
290 58.6 16994
Mar 9 290 58.6 16994
100 53.6 5360 230-100=130 53.6 6968
Mar18 150 63.6 9540 130 53.6 6968
150 63.6 9540
Mar 25 280 65.60 18368 130 53.6 6968
150 63.6 9540
280 65.60 18368
Mar 29 260 65.60 17056 130 53.6 6968
150 63.6 9540
280-260=20 65.6 1312
Total 39410 6968+9540+1312= 17820